# Sample-Efficient Reinforcement Learning from Human Feedback via Information-Directed Sampling

### v1

@@ IV-B Constructing surrogate environments as learning targets
IV-B1 A new distance measure
The discrete environment is constructed using a covering argument with suitable distance measures. Unlike previous works on standard RL settings [9, 7] that use either $\ell_{1}$-distance or KL-divergence, we propose a new distance measure between two probability measures, called the $\ell_{g}$-distance, which is better suited to our RLHF framework:
$\ell_{g}(P,Q)$ $\triangleq\sup_{o\in\mathcal{O}}\|\log P(\cdot|o)-\log Q(\cdot|o)\|_{1}$ (4)
$=\sup_{o\in\mathcal{O}}\int_{x\in\mathcal{X}}\left|\log\frac{P(x|o)}{Q(x|o)}\right|\mathrm{d}\mu_{\mathcal{X}},$
where $\mathcal{O}=\mathcal{S}\times\mathcal{A}$. To guarantee the existence of a finite coverage and that $\ell_{g}$ is well-defined, we need the following assumptions:
Assumption 1.
Assumption 2.
For any $P\in\Theta$, there exist $\beta,B>0$ such that
Note that Assumption 2 does not restrict the transition probabilities and rewards from being equal to 0; it simply assumes a lower bound for their non-zero supports.
Remark 1.
@@ IV-B1 A new distance measure
$\ell_{g}(P,Q)\triangleq\sup_{o\in\mathcal{O}}\int_{x\in\mathcal{X}}\sum_{i}\left|\log\frac{P_{i}(x|o)}{Q_{i}(x|o)}\right|\mathrm{d}\mu_{\mathcal{X}}$ (5)
where $P_{i}$ and $Q_{i}$ are the $i$-th component of $P$ and $Q$ respectively. This generalization of one-dimension case is useful for the analysis of linear RLHF problems (Theorem 3).
Remark 2.
Similar to the KL divergence, we allow for taking infinite values of $\ell_{g}$, e.g., if there exists a subset $\mathcal{X}^{\prime}\subset\mathcal{X}$ with positive measure such that $Q(x|o)=0$ but $P(x|o)$ is nonzero on $\mathcal{X}^{\prime}$, by definition we have $\ell_{g}(P,Q)=\infty$.
@@
which completes the proof of Lemma 1. ∎
@@
Definition 1 ($\epsilon$-covering number).
@@
$\forall P\in\mathcal{G},\exists P^{\prime}\in\{G_{1},...,G_{K(\mathcal{G},\epsilon)}\}:\ell_{g}(P,P^{\prime})\leq\epsilon.$ (6)
IV-B2 Partition of the environment
First, we introduce the concept of $\epsilon$-value partition, which must exist based on Assumption 1.
Definition 2 ($\epsilon$-value partition).
@@ IV-B2 Partition of the environment
$V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1})-V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}(s_{1})\leq\epsilon.$ (7)
@@
$B(\mathcal{E}_{0},\epsilon)\triangleq\{\mathcal{E}\in\Theta:\ell_{g}(\mathcal{E},\mathcal{E}_{0})\leq\epsilon\}.$
@@
$\Theta^{\epsilon}_{h,i_{h},j_{h}}\triangleq\left\{\mathcal{E}\in\Theta\mid P_{h}^{\mathcal{E}}\in B_{h}^{P}(i_{h},\delta_{P}),R_{h}^{\mathcal{E}}\in B_{h}^{R}(j_{h},\delta_{R})\right\}.$ (8)
Setting $K(\epsilon)\triangleq\prod_{h=1}^{H}K(\Theta_{h}^{P},\delta_{P})\times K(\Theta_{h}^{R},\delta_{R})$, we can then find a one-to-one mapping from $(h,i_{h},j_{h})$ to $[K(\epsilon)]$, and we obtain an $\epsilon$-value partition that satisfies $\cup_{k=1}^{K(\epsilon)}\Theta_{k}^{\epsilon}=\Theta$ [^2] If an environment $\mathcal{E}\in\Theta$ belongs to more than one partition, we will ensure it only appears in a single partition by truncating the other partitions. Now, we prove that for any $\mathcal{E},\mathcal{E}^{\prime}$ belonging to the same partition,
@@
$\leq\sum_{h=1}^{H}\mathbb{E}_{\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}\left[HB\cdot 2\delta_{P}+B\cdot 2\delta_{R}\right]=\frac{2\epsilon}{3}\leq\epsilon.$ (9)
@@
IV-B3 Construct the surrogate environment
Based on the above $\epsilon$-value partition, we explicitly construct the surrogate environment ~$\tilde{\mathcal{E}}_{t}^{*}$ for episode $t$ as:
~$\tilde{\mathcal{E}}_{t}^{*}=\tilde{\mathcal{E}}_{k,t}^{*}\text{ iff }\mathcal{E}\in\Theta_{k}^{\epsilon},$ (10)
where ~$\tilde{\mathcal{E}}_{k,t}^{*}\triangleq\mathbb{E}_{t}\left[\mathcal{E}|\mathcal{E}\in\Theta_{k}^{\epsilon}\right]$. For this surrogate environment, we have the following result.
Lemma 2.
Fix $t\in[T]$ and environment $\mathcal{E}\in\Theta$. Given the $\epsilon$-value partition $\{\Theta_{k}^{\epsilon}\}_{k=1}^{K(\epsilon)}$ (Eqn. (8)), the surrogate environment ~$\tilde{\mathcal{E}}_{t}^{*}$ constructed by Eqn. (10) satisfies the following:
@@ IV-B3 Construct the surrogate environment
    For any $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and any instance (~$(\tilde{\mathcal{E}}_{t}^{*},\mathcal{E})\sim\mathbb{P}_{t}(\tilde{\mathcal{E}}_{t}^{*},\mathcal{E})$ , it holds that
    $\ell_{g}(P_{h}^{\tilde{\mathcal{E}}_{t}^{*}},P_{h}^{\mathcal{E}})\leq\frac{\epsilon}{2BH^{2}},\quad\ell_{g}(R_{h}^{\tilde{\mathcal{E}}_{t}^{*}},R_{h}^{\mathcal{E}})\leq\frac{\epsilon}{2BH}$ (11)
@@
    $\mathbb{E}_{t}\big{[}V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})-V_{1,\pi_{\text{TS}}^{t}}^{\mathcal{E}}(s_{1}^{t})\big{]}-\mathbb{E}_{t}\big{[}V_{1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})-V_{1,\pi_{\text{TS}}^{t}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\big{]}\leq\epsilon,$ (12)
@@
(1) Note that ~$\tilde{\mathcal{E}}_{t}^{*}$ and $\mathcal{E}$ may not lie in the same partition, since $\Theta_{k}^{\epsilon}$ may not be convex under the new metric $\ell_{g}$. Nevertheless, we can use Lemma 3 to bound the $\ell_{g}$-distance between ~$\tilde{\mathcal{E}}_{t}^{*}$ and $\mathcal{E}$. Let $\mathcal{C}$ be the center of $\Theta_{k}^{\epsilon}$, we have $\ell_{g}(P_{h}^{\tilde{\mathcal{E}}_{t}^{*}},P_{h}^{\mathcal{C}})\leq 2\delta_{P}$. Then, by triangle inequality of $\ell_{g}$ (Lemma 1), we have
$\ell_{g}(P_{h}^{\tilde{\mathcal{E}}_{t}^{*}},P_{h}^{\mathcal{E}})\leq\ell_{g}(P_{h}^{\tilde{\mathcal{E}}_{t}^{*}},P_{h}^{\mathcal{C}})+\ell_{g}(P_{h}^{\mathcal{E}},P_{h}^{\mathcal{C}})\leq 3\delta_{P}=\frac{\epsilon}{2BH^{2}}.$
@@
The analysis for the reward term $\ell_{g}(R_{h}^{\tilde{\mathcal{E}}_{t}^{*}},R_{h}^{\mathcal{E}})$ is exactly the same as above, which yields the proof of the first conclusion in Lemma 2.
@@
(2) For the second property, we divide $\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})-V_{1,\pi_{\mathrm{TS}}^{t}}^{\mathcal{E}}(s_{1}^{t})\right]-\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})-V_{1,\pi_{\mathrm{TS}}^{t}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]$ into two parts.
@@
  • We first show that $\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\mathcal{E}}(s_{1}^{t})\right]=\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right].$ Let $\mathcal{E}_{t}\sim\mathbb{P}(\cdot|\mathcal{D}_{t})$ be an independent sample of $\mathcal{E}$. By the law of total expectation and the definition of ~$\tilde{\mathcal{E}}_{t}^{*}$, we have
@@
    $\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]$ $=\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\middle|\mathcal{E}\in\Theta_{k}^{\epsilon}\right]$
@@
    $=\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\tilde{\mathcal{E}}_{k,t}^{*}}(s_{1}^{t})\right].$
    Then, using the independence over layers after conditioning on $\Theta_{k}^{\epsilon}$ and the fact that $\mathcal{E}_{t}$ is independent with $\mathcal{E}$, we have
@@
    $\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]$ $=\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\int_{\mathcal{E}^{\prime}\in\Theta_{k}^{\epsilon}}\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\mathcal{E}^{\prime}}(s_{1}^{t})\right]\mathrm{d}\mathbb{P}(\mathcal{E}_{t}=\mathcal{E}^{\prime}|\mathcal{E}_{t}\in\Theta_{k}^{\epsilon})$
@@
    $=\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\int_{\mathcal{E}^{\prime}\in\Theta_{k}^{\epsilon}}\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\mathcal{E}^{\prime}}(s_{1}^{t})\middle|\mathcal{E}_{t}\in\Theta_{k}^{\epsilon}\right]\mathrm{d}\mathbb{P}(\mathcal{E}_{t}=\mathcal{E}^{\prime}|\mathcal{E}_{t}\in\Theta_{k}^{\epsilon})$
@@
    $=\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\mathcal{E}}(s_{1}^{t})\right].$
  • Next, we show that $\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})\right]-\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]\leq\epsilon.$ Adopting the same decomposition trick as in Eqn. (IV-B2), we have
    $V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})-V_{1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})$
    $\leq\sum_{h=1}^{H}\mathbb{E}_{\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}\left[HB\cdot\frac{\epsilon}{2BH^{2}}+B\cdot\frac{\epsilon}{2BH}\right]=\epsilon,$ (13)
    where the second inequality is due to Eqn. (11). Adding up the two parts yields the proof of the second property in Lemma 2. By this, we have finished the proof of Lemma 2.
It is worth noting that Eqn. (11) represents a unique property of the surrogate environment, specifically attributed to our metric $\ell_{g}$, which distinguishes it from the KL divergence. It can be proven that the $\ell_{1}$ distance also possesses this property. However, as seen in Section V, Proposition 1 cannot be guaranteed under the $\ell_{1}$ distance, making it difficult to design efficient approximation algorithms.
@@
$\rho^{P}_{t}(P)\propto\rho^{P}(P)\prod_{i=1}^{t-1}\prod_{h=1}^{H}P_{h}(s_{h+1}^{i,1}|s_{h}^{i,1},a_{h}^{i,1})$ (14)
$\rho^{R}_{t}(R)\propto\rho^{R}(R)\prod_{i=1}^{t-1}\big{(}o_{i}\sigma(r(\tau_{1}^{i})-r(\tau_{0}^{i}))+(1-o_{i})\sigma(r(\tau_{0}^{i})-r(\tau_{1}^{i}))\big{)}$ (15)
4:     Compute the surrogate environment ~$\tilde{\mathcal{E}}_{t}^{*}$, and update policy by
$\pi_{\mathrm{IDS}}^{t}=\arg\max_{\pi\in\Pi}\mathbb{E}_{t}[V_{1,\pi}^{\mathcal{E}}(s_{1})]+\frac{\lambda}{2}\mathbb{I}_{t}^{\pi}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)$

### v3

@@ 4.2 Constructing surrogate environments as learning targets
4.2.1 A new distance measure
The discrete environment is constructed using a covering argument with suitable distance measures. Unlike previous works on standard RL settings Hao and Lattimore (2022); Moradipari et al. (2023) that use either $\ell_{1}$-distance or KL-divergence, we propose a new distance measure between two probability measures, called the $\ell_{g}$-distance, which is better suited to our RLHF framework:
$\ell_{g}(P,Q)\triangleq\sup_{o\in\mathcal{O}}\|\log P(\cdot|o)-\log Q(\cdot|o)\|_{1}=\sup_{o\in\mathcal{O}}\int_{x\in\mathcal{X}}\left|\log\frac{P(x|o)}{Q(x|o)}\right|\mathrm{d}\mu_{\mathcal{X}},$ (4.2)
where $\mathcal{O}=\mathcal{S}\times\mathcal{A}$.
Remark 4.1.
@@ 4.2.1 A new distance measure
$\ell_{g}(P,Q)\triangleq\sup_{o\in\mathcal{O}}\int_{x\in\mathcal{X}}\sum_{i}\left|\log\frac{P_{i}(x|o)}{Q_{i}(x|o)}\right|\mathrm{d}\mu_{\mathcal{X}}$ (4.3)
where $P_{i}$ and $Q_{i}$ are the $i$-th component of $P$ and $Q$ respectively. This generalization of one-dimension case is useful for the analysis of linear RLHF problems (Theorem 4.16).
Remark 4.2.
To guarantee $\ell_{g}$ is well-defined, we let $\log\frac{0}{0}\triangleq 0$. Similar to the KL divergence, we allow for taking infinite values of $\ell_{g}$, e.g., if there exists a subset $\mathcal{X}^{\prime}\subset\mathcal{X}$ with positive measure such that $Q(x|o)=0$ but $P(x|o)$ is nonzero on $\mathcal{X}^{\prime}$, by definition we have $\ell_{g}(P,Q)=\infty$.
@@
which completes the proof of Lemma 4.3. ∎
To guarantee the existence of a finite coverage, we need the following assumptions:
Assumption 4.4.
Assumption 4.5.
For any $P\in\Theta$, there exist $\beta,B>0$ such that
Note that, our assumption permits the probability density to be zero, but restricts the non-zero support to have a lower bound of $\beta$. This lower bound $\beta$ is only required for Algorithm 2 in Section 5; Algorithm 1 does not rely on this assumption. In the regret upper bound of Algorithm  2, the term involving $\beta$ is $\log\frac{1}{\beta}$. Consequently, $\beta$ can be chosen to be extremely small. For instance, if $\beta=e^{-100}$ (a value far beyond the floating-point precision of modern computers), then $\log\frac{1}{\beta}=100$. Even in this case, the upper bound is only affected by a constant factor relative to the original bound.
@@
Definition 4.6 ($\epsilon$-covering number).
@@
$\forall P\in\mathcal{G},\exists P^{\prime}\in\{G_{1},...,G_{K(\mathcal{G},\epsilon)}\}:\ell_{g}(P,P^{\prime})\leq\epsilon.$ (4.4)
4.2.2 Partition of the environment
First, we introduce the concept of $\epsilon$-value partition, which must exist based on Assumption 4.4.
Definition 4.7 ($\epsilon$-value partition).
@@ 4.2.2 Partition of the environment
$V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1})-V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}(s_{1})\leq\epsilon.$ (4.5)
@@
$B(\mathcal{E}_{0},\epsilon)\triangleq\{\mathcal{E}\in\Theta:\ell_{g}(\mathcal{E},\mathcal{E}_{0})\leq\epsilon\}.$ (4.6)
@@
$\Theta^{\epsilon}_{h,i_{h},j_{h}}\triangleq\left\{\mathcal{E}\in\Theta\mid P_{h}^{\mathcal{E}}\in B_{h}^{P}(i_{h},\delta_{P}),R_{h}^{\mathcal{E}}\in B_{h}^{R}(j_{h},\delta_{R})\right\}.$ (4.7)
Setting $K(\epsilon)\triangleq\prod_{h=1}^{H}K(\Theta_{h}^{P},\delta_{P})\times K(\Theta_{h}^{R},\delta_{R})$, we can then find a bijective mapping from $(h,i_{h},j_{h})$ to $K(\epsilon)]$, and we obtain an $\epsilon$-value partition that satisfies $\cup_{k=1}^{K(\epsilon)}\Theta_{k}^{\epsilon}=\Theta$. [^2] If an environment $\mathcal{E}\in\Theta$ belongs to more than one partition, we will ensure it only appears in a single partition by truncating the other partitions. Now, we prove that for any $\mathcal{E},\mathcal{E}^{\prime}$ belonging to the same partition,
@@
$\leq\sum_{h=1}^{H}\mathbb{E}_{\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}\left[HB\cdot 2\delta_{P}+B\cdot 2\delta_{R}\right]=\frac{2\epsilon}{3}\leq\epsilon.$ (4.8)
@@
Remark 4.8.
An important distinction between the $\ell_{g}$-distance and the $\ell_{1}$-distance is that the $\epsilon$-ball under the $\ell_{g}$-distance is not convex. In other words, there exist instances of probability measures for which the $\epsilon$-ball defined in Eq. (4.6) is non-convex. We provide a counterexample in the Appendix B to demonstrate this non-convexity. While the lack of convexity does not affect the partitioning of the environment, it does influence the construction of the surrogate environment in the subsequent analysis.
4.2.3 Construct the surrogate environment
Based on the above $\epsilon$-value partition, we explicitly construct the surrogate environment $\widetilde{\mathcal{E}}_{t}^{*}$ for episode $t$ as:
$\widetilde{\mathcal{E}}_{t}^{*}=\widetilde{\mathcal{E}}_{k,t}^{*}\text{ iff }\mathcal{E}\in\Theta_{k}^{\epsilon},$ (4.9)
where $\widetilde{\mathcal{E}}_{k,t}^{*}\triangleq\mathbb{E}_{t}\left[\mathcal{E}|\mathcal{E}\in\Theta_{k}^{\epsilon}\right]$. Since $\cup_{k=1}^{K(\epsilon)}\Theta_{k}^{\epsilon}=\Theta$, for any $\mathcal{E}$, there exists $k\in[K(\epsilon)]$ such that $\mathcal{E}\in\Theta_{k}^{\epsilon}$. Hence, the surrogate environment is well defined. For this surrogate environment, we have the following result.
Lemma 4.9.
Fix $t\in[T]$ and environment $\mathcal{E}\in\Theta$. Given the $\epsilon$-value partition $\{\Theta_{k}^{\epsilon}\}_{k=1}^{K(\epsilon)}$ (Eqn. (4.7)), the surrogate environment $\widetilde{\mathcal{E}}_{t}^{*}$ constructed by Eqn. (4.9) satisfies the following:
@@ 4.2.3 Construct the surrogate environment
    For any $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and any instance $(\widetilde{\mathcal{E}}_{t}^{*},\mathcal{E})\sim\mathbb{P}_{t}(\widetilde{\mathcal{E}}_{t}^{*},\mathcal{E})$ , it holds that
    $\ell_{g}(P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}},P_{h}^{\mathcal{E}})\leq\frac{\epsilon}{2BH^{2}},\quad\ell_{g}(R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}},R_{h}^{\mathcal{E}})\leq\frac{\epsilon}{2BH}$ (4.10)
@@
    $\mathbb{E}_{t}\big{[}V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})-V_{1,\pi_{\text{TS}}^{t}}^{\mathcal{E}}(s_{1}^{t})\big{]}-\mathbb{E}_{t}\big{[}V_{1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})-V_{1,\pi_{\text{TS}}^{t}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\big{]}\leq\epsilon,$ (4.11)
@@
A necessary step in the regret analysis is to bound $\ell_{g}(P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}},P_{h}^{\mathcal{E}})$ by the radius of the $\epsilon$-balls, i.e., Eq. (4.10). Since the $\epsilon$-ball under $\ell_{1}$ is convex, the posterior mean $\widetilde{\mathcal{E}}_{k,t}^{*}$ also lies in $\Theta_{k}^{\epsilon}$. We can directly derive that $\ell_{1}(P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}},P_{h}^{\mathcal{E}})\leq 2\epsilon.$ However, $\Theta_{k}^{\epsilon}$ is not convex under the new metric $\ell_{g}$, hence $\widetilde{\mathcal{E}}_{t}^{*}$ and $\mathcal{E}$ may not lie in the same partition. Therefore, we are unable to immediately obtain Eq. (4.10). Although we cannot use the convexity of $\epsilon$-balls under $\ell_{g}$, a specific geometric property of $\ell_{g}$ (Lemma B.1) significantly simplifies our proof.
@@
(1) Based on our environment partitioning, $\Theta_{k}^{\epsilon}$ is a ball. Let $\mathcal{C}$ be the center of $\Theta_{k}^{\epsilon}$, by Lemma B.1, we have $\ell_{g}(P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}},P_{h}^{\mathcal{C}})\leq 2\delta_{P}$. Then, by triangle inequality of $\ell_{g}$ (Lemma 4.3), we have
$\ell_{g}(P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}},P_{h}^{\mathcal{E}})\leq\ell_{g}(P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}},P_{h}^{\mathcal{C}})+\ell_{g}(P_{h}^{\mathcal{E}},P_{h}^{\mathcal{C}})\leq 3\delta_{P}=\frac{\epsilon}{2BH^{2}}.$
@@
The analysis for the reward term $\ell_{g}(R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}},R_{h}^{\mathcal{E}})$ is exactly the same as above, which yields the proof of the first conclusion in Lemma 4.9.
@@
(2) For the second property, we divide $\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})-V_{1,\pi_{\mathrm{TS}}^{t}}^{\mathcal{E}}(s_{1}^{t})\right]-\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})-V_{1,\pi_{\mathrm{TS}}^{t}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]$ into two parts.
@@
  • We first show that $\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\mathcal{E}}(s_{1}^{t})\right]=\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right].$ Let $\mathcal{E}_{t}\sim\mathbb{P}(\cdot|\mathcal{D}_{t})$ be an independent sample of $\mathcal{E}$. By the law of total expectation and the definition of $\widetilde{\mathcal{E}}_{t}^{*}$, we have
@@
    $\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]$ $=\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\middle|\mathcal{E}\in\Theta_{k}^{\epsilon}\right]$
@@
    $=\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\widetilde{\mathcal{E}}_{k,t}^{*}}(s_{1}^{t})\right].$
    By the definition of $\widetilde{\mathcal{E}}_{k,t}^{*}$,
@@
    $\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\widetilde{\mathcal{E}}_{k,t}^{*}}(s_{1}^{t})\right]=\int_{\mathcal{E}^{\prime}\in\Theta_{k}^{\epsilon}}\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\mathcal{E}^{\prime}}(s_{1}^{t})\right]\mathrm{d}\mathbb{P}(\mathcal{E}_{t}=\mathcal{E}^{\prime}|\mathcal{E}_{t}\in\Theta_{k}^{\epsilon}).$
    Then, we have
@@
    $\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]$ $=\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\widetilde{\mathcal{E}}_{k,t}^{*}}(s_{1}^{t})\right]$
    $=\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\int_{\mathcal{E}^{\prime}\in\Theta_{k}^{\epsilon}}\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\mathcal{E}^{\prime}}(s_{1}^{t})\right]\mathrm{d}\mathbb{P}(\mathcal{E}_{t}=\mathcal{E}^{\prime}|\mathcal{E}_{t}\in\Theta_{k}^{\epsilon})$
    $\overset{(a)}{=}\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\int_{\mathcal{E}^{\prime}\in\Theta_{k}^{\epsilon}}\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\mathcal{E}^{\prime}}(s_{1}^{t})\middle|\mathcal{E}_{t}\in\Theta_{k}^{\epsilon}\right]\mathrm{d}\mathbb{P}(\mathcal{E}_{t}=\mathcal{E}^{\prime}|\mathcal{E}_{t}\in\Theta_{k}^{\epsilon})$
@@
    $\overset{(b)}{=}\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\mathcal{E}}(s_{1}^{t})\right].$
    where (a) uses the fact that $\pi_{\mathrm{TS}}^{t}=\pi^{*}_{\mathcal{E}_{t}^{\prime}}$, $\mathcal{E}_{t}^{\prime}$ is independent of $\mathcal{E}_{t}$, (b) follows from $\mathcal{E}_{t}$ is an independent sample of $\mathcal{E}$.
  • Next, we show that $\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})\right]-\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]\leq\epsilon.$ Adopting the same decomposition trick as in Eqn. (4.2.2), we have
    $V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})-V_{1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})$
    $\leq\sum_{h=1}^{H}\mathbb{E}_{\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}\bigg{[}HB\cdot\int_{\mathcal{S}}\left|\log\frac{P_{h}^{\mathcal{E}}(s^{\prime}|s_{h},a_{h})}{P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(s^{\prime}|s_{h},a_{h})}\right|\mathrm{d}\mu_{\mathcal{S}}+B\cdot\int_{[0,1]}\left|\log\frac{R_{h}^{\mathcal{E}}(x|s_{h},a_{h})}{R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(x|s_{h},a_{h})}\right|\mathrm{d}x\bigg{]}$
    $\leq\sum_{h=1}^{H}\mathbb{E}_{\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}\left[HB\cdot\frac{\epsilon}{2BH^{2}}+B\cdot\frac{\epsilon}{2BH}\right]=\epsilon,$ (4.12)
    where the second inequality is due to Eqn. (4.10). Adding up the two parts yields the proof of the second property in Lemma 4.9. By this, we have finished the proof of Lemma 4.9.
It is worth noting that Eqn. (4.10) represents a unique property of the surrogate environment, specifically attributed to our metric $\ell_{g}$, which distinguishes it from the KL divergence. It can be proven that the $\ell_{1}$-distance also possesses this property. However, as seen in Section 5, Proposition 5.1 cannot be guaranteed under the $\ell_{1}$-distance, making it difficult to design efficient approximation algorithms.
@@
$\rho^{P}_{t}(P)\propto\rho^{P}(P)\prod_{i=1}^{t-1}\prod_{h=1}^{H}P_{h}(s_{h+1}^{i,1}|s_{h}^{i,1},a_{h}^{i,1})$ (4.13)
$\rho^{R}_{t}(R)\propto\rho^{R}(R)\prod_{i=1}^{t-1}\big{(}o_{i}\sigma(r(\tau_{1}^{i})-r(\tau_{0}^{i}))+(1-o_{i})\sigma(r(\tau_{0}^{i})-r(\tau_{1}^{i}))\big{)}$ (4.14)
4:  Compute the surrogate environment $\widetilde{\mathcal{E}}_{t}^{*}$, and update policy by
$\pi_{\mathrm{IDS}}^{t}=\arg\max_{\pi\in\Pi}\mathbb{E}_{t}[V_{1,\pi}^{\mathcal{E}}(s_{1})]+\frac{\lambda}{2}\mathbb{I}_{t}^{\pi}\left(\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)$